OpenStudy (anonymous):

|7-2q|=3 there are 2 answers. please show work.

OpenStudy (anonymous):

7-2q=3 and 7-2q=-3

OpenStudy (anonymous):

and solve it separately

OpenStudy (anonymous):

so 7-2q=3 -7 -7 -2q= -4 -2/-2 -4/-2 q= 2

OpenStudy (anonymous):

and the other one is 7-2q=-3 -7 -7 2q=-10 2/2 -10/2 q=-5